Julie leaves home travelling at 45 mph. An hour later, her mom follows at 60 mph. How long will it take her mom to catch up to Julie?
Answer by checkley77(12844)

D=RT
D=45T FOR JULIE
D=60(T-1) FOR MOM.
45T=60(T-1)
45T=60T-60
45T-60T=-60
-15T=-60
T=-60/-15
T=4 HOURS AFTER JULIE LEAVES MOM WILL OVERTAKE HER.
PROOF:
45*4=60(4-1)
180=60*3
180=180